For the year ended December 31, 2019, Anyco Inc. reported earnings per share of $3.12. During...

  1. For the year ended December 31, 2019, Anyco Inc. reported earnings per share of $3.12.

    1. During 2020, the company had a 3-for-1 stock split. Calculate the 2019 earnings per share that will be reported in Anyco’s 2020 annual report for comparative purposes.

    2. During 2021, Anyco had an additional 2-for-1 stock split. Calculate the 2019 earnings per share that will be reported in Anyco’s 2021 annual report for comparative purposes.

    3. Assume that Anyco had issued a 10% stock dividend in 2020 and did not have a stock split. Calculate the 2019 earnings per share that will be reported in Anyco’s 2020 annual report for comparative purposes. Round your answer to two decimal places.
Answer #1

Solution 1:

2019 earnings per share that will be reported in Anyco’s 2020 annual report for comparative purposes = $3.12 / 3 = $1.04 per share

Solution 2:

2019 earnings per share that will be reported in Anyco’s 2021 annual report for comparative purposes = $3.12 / (3*2) = $0.52 per share

Solution 3:

2019 earnings per share that will be reported in Anyco’s 2020 annual report for comparative purposes = $3.12 / 1.10 = $2.84 per share